# Generative Modeling through Spectral Analysis of Koopman Operator

### v1

@@ Abstract
We propose Koopman Spectral Wasserstein Gradient Descent (KSWGD), a generative modeling framework that combines operator-theoretic spectral analysis with optimal transport. The novel insight is that the spectral structure required for accelerated Wasserstein gradient descent can be directly estimated from trajectory data via Koopman operator approximation which can eliminate the need for explicit knowledge of the target potential or neural network training. We provide rigorous convergence analysis and establish connection to Feynman-Kac theory that clarifies the method’s probabilistic foundation. Experiments across diverse settings, including compact manifold sampling, metastable multi-well systems, image generation, and high dimensional stochastic partial differential equation, demonstrate that KSWGD consistently achieves faster convergence than other existing methods while maintaining high sample quality.
Keywords Koopman operator $\cdot$ Dynamic mode decomposition $\cdot$ Spectral analysis $\cdot$ Langevin dynamics $\cdot$ Wasserstein gradient flow $\cdot$ Feynman-Kac formula $\cdot$ Generative modeling
@@ 1 Introduction
Generative modeling plays a central role in modern data science, with applications ranging from computer vision and natural language processing to scientific computing and uncertainty quantification [arjovsky2017wasserstein, baikonode, cheng2025safe, dinh2017realnvp, geng2025mean, goodfellow2014generative, ho2020denoising, nichol2021improved, radford2018improving, sohl2015deep, song2019generative, song2020score, van2016pixel, yu2025unlock]. Given samples from an unknown target distribution, the goal is to generate additional samples without explicit access to its density. Despite their empirical success, many popular generative models, such as variational autoencoders [kingma2013auto, rezende2014stochastic], generative adversarial networks [goodfellow2014generative], and diffusion or score-based models [ho2020denoising, song2019generative, song2020score, yang2023diffusion], rely on extensive neural network training, careful hyperparameter tuning, and substantial computational resources. Moreover, Langevin-type sampling and score-based approaches require access to gradients of the log-density or learned score functions, which can be costly or unstable to estimate in high-dimensional or complex settings.
@@
An alternative perspective on sampling and generative modeling is provided by Wasserstein gradient flows [ambrosio2008gradient, arbel2019maximum, fan2021variational, jordan1998variational, liu2017stein, liu2016stein, mokrov2021large], which offer a unifying variational framework for many stochastic and deterministic particle methods. Algorithms such as Stein variational gradient descent [liu2017stein, liu2016stein], affine-invariant Langevin dynamics [garbuno2020affine], and Laplacian-adjusted Wasserstein gradient descent (LAWGD) [chewi2020svgd] can be interpreted as discrete approximation of gradient flows on the space of probability measures. In particular, LAWGD achieves strong theoretical guarantees by preconditioning the gradient flow with the inverse Langevin generator $\mathcal{L}^{-1}$. However, the inverse generator is typically inaccessible due to the unknown target density.
Recent work has therefore focused on data-driven spectral approximations of the Langevin generator. A notable example is the diffusion map particle system (DMPS) [li2025diffusion], which combines diffusion maps with LAWGD to construct a kernel-based approximation of the generator from samples. Diffusion maps [coifman2006diffusion, coifman2005geometric, coifman2008diffusion, nadler2006diffusion] construct a kernel integral operator based on pairwise distances that provides a consistent approximation of the overdamped Langevin generator $\mathcal{L}$ under appropriate density normalization. This nonparametric construction is training-free and is particularly effective for distributions supported on low-dimensional manifolds. However, diffusion maps only use pairwise distances between i.i.d. samples and require careful tuning of the kernel bandwidth to approximate the generator accurately. Temporal ordering and trajectory information are not explicitly used in the operator construction, which limits the direct applicability of such methods to time-series data or prediction tasks arising from dynamical systems.
In this work, we choose an operator-theoretic perspective grounded in the Kolmogorov backward equation [engel1999one, oksendal2003stochastic], which governs the evolution of conditional expectations under stochastic dynamics. For overdamped Langevin systems, the infinitesimal generator of this evolution coincides with the Langevin generator, i.e., $\mathcal{A}=$-$\mathcal{L}$, up to a negative sign, and serves as the Koopman generator of the underlying Markov semigroup [engel1999one]. This connection allows the generator to be approximated directly from trajectory data using spectral methods such as extended dynamic mode decomposition (EDMD) [Williams_2015], kernel EDMD [kevrekidis2016kernel], and neural network-based Koopman approaches [lusch2018deep, li2017extended, xu2025reskoopnet]. While diffusion maps approximate the generator based on static pairwise geometry, Koopman-based methods can also explicitly incorporate temporal evolution through time-ordered snapshot pairs. Using Galerkin projection onto a finite-dimensional dictionary space, the spectral approximation enables the construction of the inverse operator $\mathcal{L}^{-1}$ required by LAWGD. Unlike general applications of Koopman operator theory to chaotic or complex systems, where the spectrum is non-real and eigenfunctions are non-orthogonal, the generative modeling task relies on Langevin dynamics, which satisfies the detailed balance condition. This physical property induces a crucial mathematical structure that the associated Koopman generator is self-adjoint (Hermitian) with respect to the invariant measure $\pi$. The self-adjointness guarantees that the spectrum is purely real and the eigenfunctions form an orthogonal basis of $L^{2}_{\pi}$, which is an important mathematical property and valuable to us.
Building on this insight, we propose Koopman Spectral Wasserstein Gradient Descent (KSWGD), a training-free generative modeling framework that integrates Wasserstein gradient flows with modern Koopman spectral analysis [brunton2022data, brunton2021modern, mezic2005spectral]. Replacing the inaccessible inverse Langevin generator with its Koopman-based spectral approximation leads to deterministic particle dynamics with a constant dissipation rate, preventing the vanishing-gradient behavior observed in kernel-based particle methods [duncan2023geometry, fujisawa2025on, liu2023towards, wainwright2019high]. We establish linear convergence in discrete time, together with explicit error bounds determined by spectral truncation and operator approximation accuracy. The Koopman formulation further admits a natural linear Feynman–Kac interpretation [oksendal2003stochastic], which provides a rigorous probabilistic foundation for unconditional sampling.
The main contributions of this paper are:
  1. 1.
    We proposed a training-free generative modeling framework based on Koopman spectral approximations of the inverse Langevin generator.
  2. 2.
    Convergence guarantees with explicit error bounds arising from spectral truncation and numerical approximation.
  3. 3.
    A (backward) operator-theoretic perspective that unifies Wasserstein gradient flows, Koopman theory, and generative particle systems.
  4. 4.
    Empirical validation of data in the form of both time series and static time on compact manifold, metastable systems, image data, and high-dimensional stochastic partial differential equation.
The remainder of the paper is organized as follows. Section 2 reviews the operator-theoretic background on Wasserstein gradient flows and Koopman semigroups. Section 3 introduces the proposed Koopman Spectral Wasserstein Gradient Descent (KSWGD) framework and its spectral construction. Section 4 establishes convergence guarantees and error bounds based on numerical spectral truncation. Section 5 provides a Feynman–Kac interpretation of KSWGD, clarifying its connection to unconditional sampling and potential extension to conditional sampling problems. Numerical experiments are presented in Section 6, followed by conclusion and future directions in Section 7.
2 Preliminary
@@ 2.1 Wasserstein Gradient Flow
Let $\pi(x)\propto e^{-V(x)}$ be a target distribution on $\mathbb{R}^{d}$. The Jordan–Kinderlehrer–Otto (JKO) framework [jordan1998variational] interprets evolution equations for probability measures as Wasserstein gradient flows on $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$ where $W_{2}$ denotes the $2$-Wasserstein distance [santambrogio2015optimal].
For a functional $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$, the Wasserstein gradient flow is defined through the continuity equation
$\partial_{t}\mu_{t}=\operatorname{div}(\mu_{t}\,\nabla_{W_{2}}\mathcal{F}(\mu_{t})),$
where $\nabla_{W_{2}}\mathcal{F}(\mu):\mathbb{R}^{d}\to\mathbb{R}^{d}$ denotes the Wasserstein gradient [santambrogio2015optimal]. This PDE describes the evolution of the distribution $\mu_{t}$ whose law is governed by particles moving along the velocity field
$v_{t}(x)=-\nabla_{W_{2}}\mathcal{F}(\mu_{t})(x).$
Classical examples include the Kullback–Leibler (KL) divergence $\mathrm{KL}(\mu\|\pi)\coloneqq\int\log(\mathrm{d}\mu/\mathrm{d}\pi)\,d\mu$ with Wasserstein gradient $\nabla_{W_{2}}\mathrm{KL}(\mu)=\nabla\log(\mathrm{d}\mu/\mathrm{d}\pi)$, and the chi-squared divergence $\chi^{2}(\mu\|\pi)\coloneqq\mathbb{E}_{\pi}[(\mathrm{d}\mu/\mathrm{d}\pi-1)^{2}]$ with $\nabla_{W_{2}}\chi^{2}(\mu)=2\nabla(\mathrm{d}\mu/\mathrm{d}\pi)$ [ambrosio2008gradient, santambrogio2015optimal, Tsybakov2009]. The chi-squared gradient flow enables sharper convergence analysis for deterministic particle methods [chewi2020svgd].
LAWGD. The chi-squared gradient flow $\partial_{t}\mu_{t}=2\operatorname{div}(\mu_{t}\nabla(\mathrm{d}\mu_{t}/\mathrm{d}\pi))$ admits strong theoretical guarantees but is computationally intractable, in other words, evaluating the velocity field $2\nabla(\mathrm{d}\mu_{t}/\mathrm{d}\pi)$ requires estimating the density ratio $\mathrm{d}\mu_{t}/\mathrm{d}\pi$ at each iteration, which suffers from the curse of dimensionality. To circumvent this difficulty, Laplacian-Adjusted Wasserstein Gradient Descent (LAWGD) [chewi2020svgd] applies a spectral preconditioning via the integral operator $\mathcal{K}_{\pi}$:
$\partial_{t}\mu_{t}=\operatorname{div}\!\left(\mu_{t}\,\nabla\mathcal{K}_{\pi}(\mathrm{d}\mu_{t}/\mathrm{d}\pi)\right),$ (2.1)
where $\mathcal{K}_{\pi}f(x)\coloneqq\int K(x,y)f(y)\,\mathrm{d}\pi(y)$ is the integral operator associated with a carefully chosen kernel $K$. The most important design principle here is to select $K$ such that $\mathcal{K}_{\pi}=\mathcal{L}^{-1}$ where $\mathcal{L}\coloneqq-\Delta+\langle\nabla V,\nabla\cdot\rangle$ is the Langevin generator associated with $\pi$ and the inverse is understood as the inverse of $\mathcal{L}$ restricted to the orthogonal complement of constant functions in $L^{2}_{\pi}$. Notice that, here $\mathcal{L}$ is a positive self-adjoint operator on $L^{2}_{\pi}$. For standard distribution with exponential or faster tail decay, $\mathcal{L}$ has a discrete spectrum $0=\lambda_{0}<\lambda_{1}\leq\lambda_{2}\leq\cdots$ and corresponding orthonormal eigenfunctions $\{\phi_{i}\}_{i\geq 0}$ in $L^{2}_{\pi}$ with $\phi_{0}\equiv 1$ (constant function). The LAWGD kernel is then constructed via spectral representation:
$K_{\mathcal{L}^{-1}}(x,y)\coloneqq\sum_{i=1}^{\infty}\frac{\phi_{i}(x)\phi_{i}(y)}{\lambda_{i}},$ (2.2)
where the summation excludes the zero eigenvalue to ensure well-definedness. This choice yields a remarkable dissipation identity. To see this, note that by integration by parts [chewi2020svgd, Eq. (10)]:
$\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{KL}(\mu_{t}\|\pi)=-\left\langle\nabla\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\pi},\nabla\mathcal{K}_{\pi}\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\pi}\right\rangle_{\pi}=-\left\langle\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\pi}-1,\mathcal{L}\mathcal{K}_{\pi}\left(\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\pi}-1\right)\right\rangle_{\pi}=-\chi^{2}(\mu_{t}\|\pi),$ (2.3)
where the second equality uses the Dirichlet form $\langle\nabla f,\nabla g\rangle_{\pi}=\langle f,\mathcal{L}g\rangle_{\pi}$. This constant dissipation rate implies scale-free exponential convergence $\mathrm{KL}(\mu_{t}\|\pi)\leq\mathrm{KL}(\mu_{0}\|\pi)e^{-t}$ independent of the Poincaré constant [chewi2020svgd, Theorem 4]. Particle implementation. The continuity equation (2.1) translates to particle dynamics via the velocity field $v_{t}(x)=-\nabla\mathcal{K}_{\pi}(\mathrm{d}\mu_{t}/\mathrm{d}\pi)(x)$. Using the identity
$\nabla\mathcal{K}_{\pi}f(x)=\int\nabla_{1}K(x,y)f(y)\,\mathrm{d}\pi(y)=\int\nabla_{1}K_{\mathcal{L}^{-1}}(x,y)\,\mathrm{d}\mu_{t}(y),$
where $\nabla_{1}$ denotes gradient w.r.t. the first argument, particles $\{x_{t}^{(i)}\}_{i=1}^{M}$ evolve according to
$\dot{x}_{t}^{(i)}=-\frac{1}{M}\sum_{j=1}^{M}\nabla_{1}K_{\mathcal{L}^{-1}}(x_{t}^{(i)},x_{t}^{(j)}),\;1\leq i\leq M,$ (2.4)
where $\mu_{t}\approx\frac{1}{M}\sum_{j=1}^{M}\delta_{x_{t}^{(j)}}$ is the empirical measure.
Towards data-driven implementation. The implementation of LAWGD relies on the spectral decomposition of the Langevin generator $\mathcal{L}$, which requires explicit access to the potential function $V$. In many applications, however, $V$ is unknown or computationally inaccessible, i.e., the target distribution $\pi$ may be specified only through samples, or the dynamics may arise from complex systems where no closed-form potential exists. This gap between theory and practice motivates the development of data-driven methods for spectral approximation.
@@ 2.2 Koopman Operator
Consider a dynamical system governed by the SDE
$\mathrm{d}X_{t}=b(X_{t})\mathrm{d}t+\sigma(X_{t})\mathrm{d}W_{t},$ (2.5)
where $b:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the drift, $\sigma:\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$ is the diffusion coefficient satisfying appropriate regularity condition [engel1999one], and $W_{t}$ is an $m$-dimensional Brownian motion.
Koopman operator theory [Koopman1931, koopman1932dynamical, mezic2005spectral] provides a linear framework for analyzing such systems by lifting the nonlinear dynamics to a linear operator acting on observable functions. For the continuous time stochastic system above, the Koopman operator $\mathcal{T}^{t}:L^{2}(\mathcal{X})\to L^{2}(\mathcal{X})$ acts on observable functions $g:\mathcal{X}\to\mathbb{R}$ via
$(\mathcal{T}^{t}g)(x)\coloneqq\mathbb{E}[g(X_{t})\mid X_{0}=x],$ (2.6)
where the conditional expectation is taken over all realizations of the Brownian motion given the initial state $x$. This operator is linear despite the potential nonlinearity of $b$ and $\sigma$, enabling the application of spectral methods to analyze the stochastic dynamics. The infinitesimal generator $\mathcal{A}$ of the Koopman operator is defined as
$\mathcal{A}f\coloneqq\lim_{t\to 0}\frac{\mathcal{T}^{t}f-f}{t},$
on the domain
$\mathcal{D}(\mathcal{A})=\left\{f\in L^{2}(\mathcal{X}):\lim_{t\to 0}\frac{\mathcal{T}^{t}f-f}{t}\text{ exists in }L^{2}(\mathcal{X})\right\}.$
By Itô’s formula [engel1999one], this generator has the following explicit form

### v3

# Generative Modeling through Koopman Spectral Analysis: An Operator-Theoretic Perspective

@@ Abstract
We propose Koopman Spectral Wasserstein Gradient Descent (KSWGD), a particle-based generative modeling framework that learns the Langevin generator via Koopman theory and integrates it with Wasserstein gradient descent. Our key insight is that this spectral structure of the underlying distribution can be directly estimated from trajectory data via the Koopman operator, eliminating the need for explicit knowledge of the target potential. Additionally, we prove that KSWGD maintains an approximately constant dissipation rate, thereby establishing linear convergence and overcoming the vanishing-gradient phenomenon that hinders existing kernel-based particle methods. We further provide a Feynman–Kac interpretation that clarifies the method’s probabilistic foundation. Experiments on compact manifolds, metastable multi-well systems, and high-dimensional stochastic partial differential equations demonstrate that KSWGD consistently outperforms baselines in both convergence speed and sample quality.
Machine Learning, ICML
@@ 1 Introduction
Generative modeling plays a central role in modern data science, aiming to generate new samples from an unknown target distribution without explicit access to its density. While approaches like variational autoencoders, generative adversarial networks, and diffusion models have achieved striking empirical success (Goodfellow et al., 2014; Kingma and Welling, 2013; Ho et al., 2020), they often demand extensive neural network training and delicate hyperparameter tuning. Furthermore, score-based methods rely on estimating gradients of the log-density, which can be computationally expensive or unstable in high-dimensional scientific problems.
@@
A complementary line of work formulates generative modeling as Wasserstein gradient flows on the space of probability measures (Jordan et al., 1998; Ambrosio et al., 2008). This variational perspective underpins particle-based algorithms such as Stein variational gradient descent (SVGD) (Liu and Wang, 2016) and Laplacian-adjusted Wasserstein gradient descent (LAWGD) (Chewi et al., 2020). LAWGD, in particular, achieves strong theoretical guarantees by preconditioning the gradient flow with the inverse Langevin generator $\mathcal{L}^{-1}$, thereby equalizing scales across stiff modes and accelerating convergence. However, realizing this ideal preconditioner is challenging in practice, as access to $\mathcal{L}^{-1}$ typically requires explicit knowledge of the potential function, which is unavailable in many generative settings.
Recent data-driven approaches, such as the diffusion map particle system (DMPS) (Li and Marzouk, 2025), approximate the generator from samples via diffusion maps. While training-free, this method rely on static pairwise distances between i.i.d. samples and often require careful bandwidth tuning. Crucially, this approach does not explicitly leverage temporal trajectory information, and its kernel-based spectral approximation does not scale well to high-dimensional settings.
In this work, we introduce an operator-theoretic perspective that bridges this gap. We leverage the fundamental connection that the Langevin generator $\mathcal{L}$ coincides (up to a sign) with the infinitesimal generator of the Koopman operator associated with the overdamped Langevin dynamics. This insight allows us to approximate the spectrum of $\mathcal{L}$ directly from trajectory data using Koopman spectral methods (Williams et al., 2015; Kevrekidis et al., 2016), rather than relying on static geometry. By truncating this data-driven spectral decomposition, we construct a finite-rank approximation of $\mathcal{L}^{-1}$ that is naturally informed by the system’s temporal evolution.
Based on this connection, we propose Koopman Spectral Wasserstein Gradient Descent (KSWGD), a particle-based framework that integrates Wasserstein gradient flows with modern Koopman theory developed by Mezić (2005, 2021). KSWGD replaces the inaccessible preconditioner in LAWGD with a Koopman-based spectral approximation. This leads to deterministic particle dynamics with an approximately constant dissipation rate, effectively mitigating the vanishing-gradient issues plaguing existing kernel methods. The resulting algorithm is training-free once the Koopman basis is estimated and accommodates both time-series and static data.
Our contributions are threefold. First, we provide a rigorous convergence analysis, establishing linear convergence in discrete time, with explicit error bounds separating spectral truncation from estimation error. Second, we establish a connection to Feynman–Kac theory (Øksendal, 2003), clarifying the method’s probabilistic foundation. Finally, empirical results on compact manifolds, metastable multi-well systems, and high-dimensional stochastic partial differential equations demonstrate that KSWGD consistently achieves faster convergence than existing particle methods while maintaining high sample quality.
2 Preliminaries
This section establishes the theoretical foundations for our method: the geometry of Wasserstein gradient flows in Section 2.1 and the spectral properties of the Koopman operator in Section 2.2.
@@ 2.1 Wasserstein Gradient Flow
Let $\pi(x)\propto e^{-V(x)}$ be a target density on $\mathbb{R}^{d}$. The Jordan–Kinderlehrer–Otto (JKO) framework (Jordan et al., 1998; Otto, 2001) interprets the evolution of probability measures as a gradient flow in the Wasserstein space $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$, where $W_{2}$ denotes the $2$-Wasserstein distance. The flow follows the continuity equation
$\partial_{t}\mu_{t}+\operatorname{div}(\mu_{t}\,v_{t})=0$
driven by the velocity field (Santambrogio, 2015)
$v_{t}=-\nabla_{W_{2}}\mathcal{F}(\mu_{t}).$
Let $\mu\ll\pi$ and write the density ratio $\rho\coloneqq\frac{\mathrm{d}\mu}{\mathrm{d}\pi}.$ We then focus on the $\chi^{2}$-divergence $\chi^{2}(\mu\|\pi)\coloneqq\int(\rho-1)^{2}\,\mathrm{d}\pi$, which provides sharper convergence guarantees than the KL divergence (Chewi et al., 2020). Its Wasserstein gradient is
$\nabla_{W_{2}}\chi^{2}(\mu_{t})=2\nabla\rho_{t},$
leading to the flow
$\partial_{t}\mu_{t}=2\operatorname{div}(\mu_{t}\nabla\rho_{t}).$
Direct simulation of this flow is computationally intractable due to the difficulty of estimating the density ratio $\rho_{t}=\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\pi}$ and its gradients in high dimensions.
To overcome this, we employ spectral preconditioning via an integral operator $\mathcal{K}_{\pi}$. Ideally, setting $\mathcal{K}_{\pi}$ to the inverse Langevin generator $\mathcal{L}^{-1}$ acts as a Newton-type update, equalizing spectral scales. The preconditioned flow is given by:
$\partial_{t}\mu_{t}=\operatorname{div}\!\left(\mu_{t}\,\nabla\mathcal{K}_{\pi}\rho_{t}\right),$ (1)
where $(\mathcal{K}_{\pi}f)(x)\coloneqq\int K(x,y)f(y)\mathrm{d}\pi(y)$ for a kernel $K$. While the ideal choice $K_{\mathcal{L}^{-1}}$ is typically unknown, our goal in Section 3 is to approximate it purely from data.
Particle implementation.
The particle dynamics are derived from the velocity field
$v_{t}(\cdot)=-\nabla\mathcal{K}_{\pi}\rho_{t}(\cdot).$
Crucially, the dependence on the density ratio $\rho_{t}$ can be eliminated using the identity $\rho_{t}(y)\mathrm{d}\pi(y)=\mathrm{d}\mu_{t}(y)$:
$\nabla\mathcal{K}_{\pi}\rho_{t}(x)$ $=\int_{\mathbb{R}^{d}}\nabla_{1}K(x,y)\,\rho_{t}(y)\,\mathrm{d}\pi(y)$
$=\int_{\mathbb{R}^{d}}\nabla_{1}K(x,y)\,\mathrm{d}\mu_{t}(y).$
This allows us to evaluate the gradient solely using samples from $\mu_{t}$. Approximating $\mu_{t}\approx\frac{1}{M}\sum_{j=1}^{M}\delta_{x_{t}^{(j)}}$, we obtain the computable particle update:
$\dot{x}_{t}^{(i)}=-\frac{1}{M}\sum_{j=1}^{M}\nabla_{1}K_{\mathcal{L}^{-1}}\!\left(x_{t}^{(i)},x_{t}^{(j)}\right),\quad 1\leq i\leq M.$ (2)
In Section 3, we will construct a data-driven approximation of $K_{\mathcal{L}^{-1}}$ using Koopman spectral analysis.
@@ 2.2 Koopman Operator
Consider a dynamical system on $\mathbb{R}^{d}$ governed by the SDE
$\mathrm{d}X_{t}=b(X_{t})\mathrm{d}t+\sigma(X_{t})\mathrm{d}W_{t},$ (3)
where $b:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the drift, $\sigma:\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$ is the diffusion coefficient, and $W_{t}$ is a standard $m$-dimensional Brownian motion. Koopman operator theory (Koopman, 1931; Koopman and Neumann, 1932; Mezić, 2005) lifts this nonlinear dynamics to a linear operator $\mathcal{T}^{t}$ acting on observable functions $g\in L^{2}(\pi)$ via conditional expectations:
$(\mathcal{T}^{t}g)(x)\coloneqq\mathbb{E}[g(X_{t})\mid X_{0}=x].$ (4)
Crucially, $\mathcal{T}^{t}$ is linear even if the underlying SDE is nonlinear, enabling the use of spectral methods. The evolution of observables is governed by the infinitesimal generator $\mathcal{A}$, defined as
$\mathcal{A}f\coloneqq\lim_{t\to 0}\frac{\mathcal{T}^{t}f-f}{t}$
on its domain $\mathcal{D}(\mathcal{A})\subset L^{2}(\pi)$. By Itô’s lemma (Engel et al., 1999), this generator admits the explicit second-order differential form: